solve the following equation:
x+1=sqrt7x-5

Answer by mananth(16946)

x+1=sqrt7x-5
square both sides of the equation
(x+1)^2=(sqrt7x-5)^2
x^2+2x+1 = 7x-5
x^2-5x+6=0
x^2-3x-2x+6=0
x(x-3)-2(x-3)=0
(x-3)(x-2)=0
x=3, or 2